### v1

@@ 3.1 Luminosities
| NGC 3293 | This work | Dufton et al. (2006) | |||||||||
| ESL No. | RV | A(Si) | A(O) | ||||||||
| 002 | 14 | 65 | 42 | 23000 | 2.8 | 100 | |||||
| 003 | 12 | 54 | 27 | 21100 | 2.9 | 80 | |||||
| 004 | 15 | 95 | 30 | mean | mean | 21500 | 3.0 | 105 | |||
| 005 | 12 | 180 | 15 | mean | mean | 21500 | 3.0 | 195 | |||
| 006 | 15 | 200 | 15 | mean | mean | 21500 | 3.15 | 200 | |||
| 007 | 12 | 28 | 26 | 22800 | 3.1 | 65 | |||||
| 008 | 12 | 130 | 30 | mean | mean | 21500 | 3.25 | 140 | |||
| 010 | 11 | 35 | 30 | 21000 | 3.15 | 70 | |||||
| 012 | 12 | 90 | 15 | mean | mean | 21500 | 3.3 | 100 | |||
| 015 | 6 | 260 | 15 | mean | mean | 25000 | 3.9 | 260 | |||
| 016 | 10 | 55 | 15 | mean | mean | 19200 | 4.0 | 110 | |||
| 019 | 8 | 120 | 15 | mean | mean | 25000 | 3.85 | 120 | |||
| 023 | 5 | 170 | 15 | mean | mean | 20500 | 3.4 | 160 | |||
| 024 | 5 | 160 | 15 | mean | mean | 20500 | 3.5 | 135 | |||
| 025 | 5 | 215 | 15 | mean | mean | 21100 | 3.7 | 215 | |||
| 028 | 4 | 200 | 15 | mean | mean | 19400 | 3.65 | 215 | |||
| 030 | 6 | 220 | 15 | mean | mean | 19800 | 3.7 | 205 | |||
| 031 | 6 | 230 | 15 | mean | mean | 17400 | 4.3 | 230 | |||
| 038 | 6 | 235 | 15 | mean | mean | 20600 | 3.95 | 235 | |||
@@
Given the brightness of these stars, it is to be expected that uncertainties in the apparent magnitudes are dominated by systematics and possible variability. While Johnson photometry with notional 1 errors of 0.03 has been published by Evans et al. (2005) and Baume et al. (2003), there are some clear outliers among the sources in common, with some differences ranging up to approximately 3 in both and . We therefore compared these datasets with the photometry derived from XP spectra, that exists for 17 of our sources, and with the photometry derived from the 4 sources with HST spectrophotometry discussed above. We followed the procedure outlined in Maíz Apellániz (2007), and used the zero-points noted therein. Excellent agreement was found between the and HST results, with and . However the magnitudes are consistently brighter and bluer than published data by up to 0.1 in and . We suspect that some of these differences may be attributed to variability (many sources are Cepheid variables), and the published data are single epoch measurements, while the spectra were taken over multiple epochs. The photometry derived from the XP spectra is therefore adopted in the present work, with uncertainties of 0.02 being used for both and . Stars ESL 028 and 030 do not have XP spectra, for these we used transformation formulae (Riello et al., 2021; Bessell, 2000) to convert native photometry to the and , adopting slightly larger uncertainties of 0.03 for these values.
@@
Traditional isochrone fitting done using non-rotating models has yielded a range of ages for this cluster. E.g., Baume et al. (2003) estimated Myr from the apparent nuclear age of the turnoff and Myr from the luminosity of the pre-main sequence stars relative to the zero age main-sequence based on a significantly larger distance estimate of 2750 pc, while more recently Bisht et al. (2021) found an age of Myr. Riello et al. (2021) included estimates of the effects of rotational distortion and mixing and suggested a significantly older age near 20 Myr. However, this result will be sensitive to how the adopted details of the interior mixing affect stellar lifetime, and still assumed that most observed stars in this cluster could be explained as products of single star evolution.
@@
In NGC 3293, most of the stars considered by Proffitt et al. (2016), as well as the brighter ones in this paper may well correspond to eMSTO stars, and as such may well be the products of binary interaction. The fainter stars, ESL 23, 24, 28, 30, 31, and 38, would in this interpretation be part of the red main-sequence due to their fast rotation and can be treated as stars that have evolved as single rotating stars without effects from mergers or binary interactions. The eMSTO stars might be expected to be depleted in boron, especially if these are mass transfer remnants, while the red sequence stars should show more modest boron depletion due to the effects of rotationally driven mixing.
@@ 4 Fitting of UV Spectra
To estimate the errors, in addition to calculating the formal fitting errors, we perturb a number of our assumptions and then redetermine the best fit boron abundance for each case. Our detailed procedure is as follows:
@@
| 2066.430 | 2065.771 | B III | -0.413 | 0.000 | Mean of hyperfine, scaled by 0.8 |
| 2066.115 | 2065.455 | B III | -1.014 | 0.000 | Mean of hyperfine, scaled by 0.2 |
@@
Even in the absence of significant gravity darkening, there are other uncertainties to consider. For the relatively narrow-lined star ESL 07 and ESL 016, the Cr ii ISM feature obscures most of the equivalent width of the stellar boron feature. While in principle one could attempt to fit a model of the ISM to the three members of the Cr ii triplet that are covered by our data, in practice the COS G185M does not quite resolve the individual components, the strength of which appear to vary significantly and independently between site-lines in this cluster. In addition, the other two components of the triplet fall in regions where the stellar opacity is changing rapidly introducing additional uncertainties. Given other uncertainties in the spectral syntheses, in practice we have only adopted upper limits for the boron abundance in these two stars. This also provides an illustration of the limits to which we can trust our spectral synthesis. In this star, with our default line list, the blend just red-ward of the boron feature in ESL 016, (Fig. 3a), includes a bit too much absorption even in the absence of any boron. Reducing the Mn iii 2066.546 Å line by 0.35 dex and of the Ni iii 2066.824 Å line by 0.3 dex (restoring the latter to its canonical value), resolves that discrepancy, (Fig. 3b), allowing a larger boron abundance. So for this star, and the similarly narrow-lined ESL 07, after consideration of such uncertainties, we adopt only upper limits to the boron abundance.
@@
For the lower mass stars in our sample, ESL 23, 24, 28, 30, 31, and 38, we find significantly larger boron abundances, and applying the same line list change we illustrated for ESL 016 would only increase the abundances by 0.1 to 0.15 dex. Our fits for ESL 28, 30, 31, and 38 are shown in Figure 4, and our new fits to ESL 023 and 024 are qualitatively similar to those of Proffitt et al. (2016).
@@
The more evolved broad-lined stars are subject to additional uncertainties, due both to their apparently lower boron abundances and to the larger impact expected from rotational distortions. For ESL 005, (Fig. 5a), there does appear to be some additional opacity at the location of the boron line, but the overall shape of this spectral region is not well matched by the synthesis. For ESL006, (Fig. 5b), the overall fit is somewhat better, but the best fit boron abundance can not be distinguished from zero. Both stars show extra absorption in the blue wing of the strong blend centered around the Ti iv 2066.241 Å and Fe iii 2066.915Å lines. In ESL 015, (Fig. 5c), this extra absorption near 2067 Å is even more prominent and we have excluded the affected region from our fit to the boron feature.
@@
Hunter et al. (2009) had measured nitrogen abundances for most stars in their NGC 3293 sample, finding that, apart from ESL 01 and 02, there was otherwise no evidence for enhanced nitrogen in any of their stars. However, they did not derive an abundance for ESL 015, and in our initial attempt to do this, we were surprised when a preliminary fit to the 3995 Å N ii feature using the parameters in Table 2 suggested a nitrogen enhancement of about 0.3 dex relative to the other stars, especially given our apparent detection of a large boron abundance for this star. However, a closer examination of the optical spectrum suggests that many of the lines are affected by additional narrow absorption features that are not reproduced by the broad-lined synthetic spectrum. This can be seen, for example, in the Si iii triplet near 4560 Å (see 6), where there appear to be additional sharp features shifted with respect to the best fit synthetic spectrum by approximately , , and km/s, with the same features appearing to repeat in each member of the triplet. Such line profile variations have been seen in other Cephei stars, (e.g. Telting & Schrijvers (1998); Schrijvers et al. (2004). Similar effects in the weaker and noisier N ii line could easily cause an overestimate of that line’s equivalent width. Until the impact of such line profile variations on abundance determinations can be taken into account, results for this star in particular should be treated with caution. While a number of our other targets are Cephei variables, none of these appear to show such obvious distortions of the line profiles.
@@ 5 Comparison with Stellar Model Predictions
Since rotationally induced mixing has been suggested as depletion mechanism, (Fliegner et al., 1996; Venn et al., 1996; Frischknecht et al., 2010; Langer et al., 2010; Brott et al., 2011a), we computed a dense grid of rotating single star models with MESA, (Jin et al., 2023), with most of the physics assumptions as in the work of Brott et al. (2011a). We use the canonical rotational mixing efficiency based on a theoretical work by Chaboyer & Zahn (1992) that is widely adopted in the literature (, cf. Heger et al., 2000; Marchant, 2017; Menon et al., 2021; Keszthelyi et al., 2022). Note that Brott et al. (2011a) adopted slightly weaker rotational mixing efficiency than the canonical value . We then used these models to produce synthetic populations for comparison with our observational data. These predictions are based on a population synthesis assuming a Salpeter (1955) initial mass function, a distribution of initial rotational velocities for early B dwarfs in the Tarantula region of the LMC given by Dufton et al. (2013), and an initial boron abundance of 2.76, (matching the initial estimated solar abundance of Asplund et al. (2021)).
@@
The adopted initial rotational velocity distribution is bimodal, with about one quarter of the stars being slow rotators with a distribution peak at , and three quarters are fast rotators with a distribution peaking at , (cf. tab. 6 of Dufton et al., 2013). Similar bimodal rotational velocity distributions are also found in young star clusters, where they manifest as split main sequences in high precision color-magnitude diagrams, (e.g., Milone et al., 2023; Wang et al., 2023), and where they are also spectroscopically verified, (Kamann et al., 2023; Bodensteiner et al., 2023). Notably, the origin of the bimodality is still under debate, (cf. Bastian et al., 2020; Wang et al., 2022), and the boron surface abundance distribution may hold interesting clues.
@@
To test the concept of boron depletion by rotational mixing, a Hunter-type diagram (surface abundance versus ) appears well suited. It has been employed repeatedly using the nitrogen surface abundance, (Hunter et al., 2008a, 2009; Brott et al., 2011a; Mahy et al., 2015; Grin et al., 2017), as well as boron, (Frischknecht et al., 2010; Langer et al., 2010). We show in Fig. 7 the boron Hunter diagram of all main sequence stars in NGC 3293 for which boron abundances have been measured or constrained so far. Since faster rotation is thought to lead to faster boron depletion, a corresponding trend may be expected, but the bulk of the data does not show such a trend. However, a trend of a decreasing boron abundance for stars with larger v sin i is expected for an unbiased massive single star population produced via constant star formation, even though with considerable scatter, (Langer et al., 2010). Our sample deviates from such a population. Instead of sampling random ages, our targets belong to an open star cluster, and can be therefore considered as coeval. As such, they contain stars near the cluster turn-off, which are close to ending core hydrogen burning, and stars of roughly half the turn-off mass and bolometrically 30-times fainter (cf., Fig.2), which so far consumed just less than 10% of their initial hydrogen in their core. Stellar models including rotational mixing do predict a faster boron depletion in more rapidly rotating stars. In addition, the depletion timescale drops rapidly with stellar mass, such that in two coeval stellar models with the same rotational velocity, the more massive one is expected to show more boron depletion, (cf., fig. 11 of Brott et al., 2011a).
@@
A closer look shows, however, that the slowest rotators in our sample are all strongly boron depleted. In particular for the mid- and high-luminosity groups, our models can not reproduce all observed stars with the same mixing efficiency and age. We consider this as an indication that a different boron depletion mechanism operated in the slow rotators, and postulate here, that the five stars with the smallest v sin i (ESL 16, 7, 10, 3 and 2) are intrinsic slow rotators, i.e., belong to the 25% of stars with a peculiar formation history, which may not be reflected by our single star models. While for each individual star this can not be proven, the likelihood that, e.g., ESL 16 belongs to the population of rapid rotators is less than 20%.
The remaining stars do show both expected trends, a stronger boron depletion for faster rotation and for higher luminosity, as predicted by our models, which can be exploited quantitatively. A close inspection of panels a-c of Fig. 7 shows that in all three luminosity ranges, less mixing than adopted in the stellar models would lead to a better agreement. This is even more so when an older age of NGC 3293 is considered, as shown by Panels d and e. Panel f in Fig. 7 illustrates the effect of reducing the efficiency parameter for rotational mixing by 50%, (). Notably, an older age can be compensated by a smaller mixing efficiency. Since an age of 12 Myr is at the lower end of the most recently proposed ages for NGC3293, (cf., Morel et al., 2022), we conclude that rotational mixing is likely less efficient than assumed in our models. Assuming we do interpret the small additional dip in the spectrum of ESL 05 (5a, relative to the otherwise similar ESL 06 (5b) as a detection of boron, then the resulting boron abundance of this star may provide the strongest constraint on a reduced rotational mixing efficiency.
@@
Perhaps, ESL 2 in our sample, and to a lesser extent also ESL 3, 7, can be considered as a smoking gun in this context. ESL 2 is an apparently slow rotator, strongly boron depleted, and the only star in our sample which is nitrogen-enriched. Its location in the HR diagram clearly identifies it as a blue straggler (Fig. 2). Blue stragglers are thought to be the products of binary mass transfer or stellar mergers, (Schneider et al., 2015). According to recent models, merger products in the mass range considered here may be slow rotators (Schneider et al., 2019) and show various degrees of nitrogen enrichments, from none to strong, (Glebbeek et al., 2013). Wang et al. (2022) have proposed that the slow rotators which form the bluer main sequence in young star clusters originate from stellar mergers. The boron abundances found in our sample stars may therefore fit into a consistent picture, with the majority of fast rotators enriching nitrogen and depleting boron by rotational mixing, and the slow rotators having a history of strong binary interaction. Slow rotators in which boron is not fully destroyed could support the suggestion of pre-main sequence mergers, (Tokovinin & Moe, 2020; Wang et al., 2022). To turn this idea from a working hypothesis into a theory, significantly more boron measurements, as well as merger models that include proton captures on boron, are required.

### v2

@@ 3.3 Luminosities
Given the brightness of these stars, it is to be expected that uncertainties in the apparent magnitudes are dominated by systematics and possible variability. While Johnson photometry with notional 1 errors of 0.03m has been published by Evans et al. (2005) and Baume et al. (2003), there are some clear outliers among the sources in common, with some differences ranging up to approximately 3 in both and . We therefore compared these datasets with the photometry derived from XP spectra, that exists for 17 of our sources, and with the photometry derived from the 4 sources with HST spectrophotometry discussed above. We followed the procedure outlined in Maíz Apellániz (2007), and used the zero-points noted therein. Excellent agreement was found between the and HST results, with and . However the magnitudes are consistently brighter and bluer than published data by up to 0.1m in and . We suspect that some of these differences may be attributed to variability (many sources are Cepheid variables), and the published data are single epoch measurements, while the spectra were taken over multiple epochs. The photometry derived from the XP spectra is therefore adopted in the present work, with uncertainties of 0.02m being used for both and . Stars ESL 028 and 030 do not have XP spectra, for these we used transformation formulae (Riello et al., 2021; Bessell, 2000) to convert native photometry to the and , adopting slightly larger uncertainties of 0.03m for these values.
@@
Traditional isochrone fitting done using non-rotating models has yielded a range of ages for this cluster. E.g., Baume et al. (2003) estimated Myr from the apparent nuclear age of the turnoff and Myr from the luminosity of the pre-main sequence stars relative to the zero age main-sequence based on a significantly larger distance estimate of 2750 pc, while more recently Bisht et al. (2021) found an age of Myr. Morel et al. (2022) included estimates of the effects of rotational distortion and mixing and suggested a significantly older age near 20 Myr. However, this result will be sensitive to how the adopted details of the interior mixing affect stellar lifetime, and still assumed that most observed stars in this cluster could be explained as products of single star evolution.
@@
In NGC 3293, most of the stars considered by Proffitt et al. (2016), as well as the brighter ones in this paper may well correspond to eMSTO stars, and as such may well be the products of binary interaction. The fainter stars, ESL 23, 24, 28, 30, 31, and 38, would in this interpretation be part of the red main-sequence due to their fast rotation and can be treated as stars that have evolved as single rotating stars without effects from mergers or binary interactions. The eMSTO stars might be expected to be depleted in boron, especially if these are mass transfer remnants (see fig. 9 in Langer, 2012; Wellstein et al., 2001), while the red sequence stars should show more modest boron depletion due to the effects of rotationally driven mixing.
@@ 4 Fitting of UV Spectra
As the B III feature is heavily blended with lines of iron group elements, before fitting this line, it is essential to obtain the best possible fit to the line profiles of the iron group elements. In practice, this requires fine tuning the micro-turbulence, and, especially for the lower mass stars in our sample, the value of implied by the UV spectra is often substantially lower than the value determined from the optical data. However, if this lower is actually compensating for uncertainties in the non-LTE model of the iron ionization equilibrium, then one might argue that the higher, optically determined values are more appropriate for the synthesis of the boron feature itself. While the need to fit the overall UV spectrum precludes us from getting a useful overall fit with the higher value, we can easily determine the change in abundance needed to yield the same equivalent width for the boron feature with the alternate micro-turbulence. Even for ESL 38, where the optically determined value of km/s is much larger than the km/s needed for the UV, the implied change in the boron abundances is only dex. For ESL 19, 23, 24, 30, and 31 the differences range from to dex with a mean difference of about dex. For the boron feature, the effect of vmicro uncertainties is somewhat muted not only because these lines are relatively weak, but also because the thermal broadening of the boron line is relatively large (at 20000K, km/s for boron vs km/s for iron), and the boron line is additionally split by the isotopic and hyperfine structure, which further reduces the saturation at small values of .
To estimate other contributions to the final errors, we perturb a number of our assumptions and then redetermine the best fit boron abundance for each case. Our detailed procedure is as follows:
@@
| 2066.430 | 2065.771 | 11B III | -0.413 | 0.000 | Mean of hyperfine, scaled by 0.8 |
| 2066.115 | 2065.455 | 10B III | -1.014 | 0.000 | Mean of hyperfine, scaled by 0.2 |
@@
Even in the absence of significant gravity darkening, there are other uncertainties to consider. For the relatively narrow-lined star ESL 07 and ESL 16, the Cr II ISM feature obscures most of the equivalent width of the stellar boron feature. While in principle one could attempt to fit a model of the ISM to the three members of the Cr II triplet that are covered by our data, in practice the COS G185M does not quite resolve the individual components, the strength of which appear to vary significantly and independently between site-lines in this cluster. In addition, the other two components of the triplet fall in regions where the stellar opacity is changing rapidly introducing additional uncertainties. Given other uncertainties in the spectral syntheses, in practice we have only adopted upper limits for the boron abundance in these two stars. This also provides an illustration of the limits to which we can trust our spectral synthesis. In this star, with our default line list, the blend just red-ward of the boron feature in ESL 16, (Fig. 3a), includes a bit too much absorption even in the absence of any boron. Reducing the Mn III 2066.546 Å line by 0.35 dex and of the Ni III 2066.824 Å line by 0.3 dex (restoring the latter to its canonical value), resolves that discrepancy, (Fig. 3b), allowing a larger boron abundance. So for this star, and the similarly narrow-lined ESL 07, after consideration of such uncertainties, we adopt only upper limits to the boron abundance.
@@
For the lower mass stars in our sample, ESL 23, 24, 28, 30, 31, and 38, we find significantly larger boron abundances, and applying the same line list change we illustrated for ESL 16 would only increase the abundances by 0.1 to 0.15 dex. Our fits for ESL 28, 30, 31, and 38 are shown in Figure 4, and our new fits to ESL 23 and 24 are qualitatively similar to those of Proffitt et al. (2016).
@@
The more evolved broad-lined stars are subject to additional uncertainties, due both to their apparently lower boron abundances and to the larger impact expected from rotational distortions. For ESL 05, (Fig. 5a), there does appear to be some additional opacity at the location of the boron line, but the overall shape of this spectral region is not well matched by the synthesis. For ESL006, (Fig. 5b), the overall fit is somewhat better, but the best fit boron abundance can not be distinguished from zero. Both stars show extra absorption in the blue wing of the strong blend centered around the Ti IV 2066.241 Å and Fe III 2066.915Å lines. In ESL 15, (Fig. 5c), this extra absorption near 2067 Å is even more prominent and we have excluded the affected region from our fit to the boron feature.
Hunter et al. (2009) had measured nitrogen abundances for most stars in their NGC 3293 sample, finding that, apart from ESL 01 and 02, there was otherwise no evidence for enhanced nitrogen in any of their stars. Morel et al. (2022) finds slightly higher mean abundances in NGC 3293 than did Hunter for both nitrogen (7.72 vs 7.60 respectively) and carbon (8.13 vs 7.97). We list the carbon and nitrogen abundances they found for our individual program stars in Table 2. While for the most part the two sets of results show similar trends, for a few individual stars Morel et al. (2022) finds nitrogen abundances that appear to be slightly enhanced. The departure is most significant for ESL 30, where Morel found an abundance of vs Hunter’s result of . However, Morel also finds a similarly enhanced carbon abundance for this star and so the carbon to nitrogen ratio is still close to the mean. Less significant apparent enhancements are also found by Morel for ESL 23 and ESL 24. However, it should also be remembered that for all four of these stars, Morel adopted a fixed value of that is significantly smaller than the value determined by Hunter, or by us above in section 3.2, and this will impact the derived abundance.
@@
However, neither Hunter et al. (2009) nor Morel et al. (2022) derived a nitrogen abundance for ESL 15, and in our initial attempt to do this, we were surprised when a preliminary fit to the 3995 Å N II feature using the parameters in Table 3 suggested a nitrogen enhancement of about 0.3 dex relative to the other stars, especially given our apparent detection of a large boron abundance for this star. However, a closer examination of the optical spectrum suggests that many of the lines are affected by additional narrow absorption features that are not reproduced by the broad-lined synthetic spectrum. This can be seen, for example, in the Si III triplet near 4560 Å (see 6), where there appear to be additional sharp features shifted with respect to the best fit synthetic spectrum by approximately , , and km/s, with the same features appearing to repeat in each member of the triplet. Such line profile variations have been seen in other Cephei stars, (e.g. Telting & Schrijvers (1998); Schrijvers et al. (2004). Similar effects in the weaker and noisier N II line could easily cause an overestimate of that line’s equivalent width. Until the impact of such line profile variations on abundance determinations can be taken into account, results for this star in particular should be treated with caution. While a number of our other targets are Cephei variables, none of these appear to show such obvious distortions of the line profiles.
Based on our current evaluation of the data, we conclude that there is no convincing evidence for significant nitrogen enhancement in any of our program stars apart from ESL 02 and possibly ESL 15. However, a more detailed reexamination of all of the available nitrogen lines to set tighter limits, including a more consistent understanding of micro-turbulence, would be worthwhile.
@@ 5 Comparison with Stellar Model Predictions
Since rotationally induced mixing has been suggested as depletion mechanism, (Fliegner et al., 1996; Venn et al., 1996; Frischknecht et al., 2010; Langer et al., 2010; Brott et al., 2011a), we computed a dense grid of rotating single star models with MESA, (Jin et al., 2023), with most of the physics assumptions as in the work of Brott et al. (2011a). For our initial calculations, we used the canonical rotational mixing efficiency based on a theoretical work by Chaboyer & Zahn (1992) that is widely adopted in the literature, (, cf. Heger et al., 2000; Marchant, 2017; Menon et al., 2021; Keszthelyi et al., 2022). Note that Brott et al. (2011a) adopted , a value slightly weaker than the canonical value. We then used these models to produce synthetic populations for comparison with our observational data. These predictions are based on a population synthesis assuming a Salpeter (1955) initial mass function, a distribution of initial rotational velocities for early B dwarfs in the Tarantula region of the LMC given by Dufton et al. (2013), and an initial boron abundance of 2.76, (matching the initial estimated solar abundance of Asplund et al. (2021)).
@@
The adopted initial rotational velocity distribution is bimodal, with about one quarter of the stars being slow rotators with a distribution peak at km s-1, and three quarters are fast rotators with a distribution peaking at km s-1, (see tab. 6 of Dufton et al., 2013). Similar bimodal rotational velocity distributions are also found in young star clusters, where they manifest as split main sequences in high precision color-magnitude diagrams, (e.g., Milone et al., 2023; Wang et al., 2023), and where they are also spectroscopically verified, (Kamann et al., 2023; Bodensteiner et al., 2023). Notably, the origin of the bimodality is still under debate, (cf. Bastian et al., 2020; Wang et al., 2022), and the boron surface abundance distribution may hold interesting clues.
@@
To test the concept of boron depletion by rotational mixing, a Hunter-type diagram (surface abundance versus ) appears well suited. It has been employed repeatedly using the nitrogen surface abundance, (Hunter et al., 2008a, 2009; Brott et al., 2011a; Mahy et al., 2015; Grin et al., 2017), as well as boron, (Frischknecht et al., 2010; Langer et al., 2010). We show in Fig. 7 the boron Hunter diagram of all main sequence stars in NGC 3293 for which boron abundances have been measured or constrained so far. Since faster rotation is thought to lead to faster boron depletion, a corresponding trend may be expected, but the bulk of the data does not show such a trend. However, a trend of a decreasing boron abundance for stars with larger is expected for an unbiased massive single star population produced via constant star formation, even though with considerable scatter, (Langer et al., 2010). Our sample deviates from such a population. Instead of sampling random ages, our targets belong to an open star cluster, and can be therefore considered as coeval. As such, they contain stars near the cluster turn-off, which are close to ending core hydrogen burning, and stars of roughly half the turn-off mass and bolometrically 30-times fainter (cf., Fig.2), which so far consumed just less than 10% of their initial hydrogen in their core. Stellar models including rotational mixing do predict a faster boron depletion in more rapidly rotating stars. In addition, the depletion timescale drops rapidly with stellar mass, such that in two coeval stellar models with the same rotational velocity, the more massive one is expected to show more boron depletion, (cf., fig. 11 of Brott et al., 2011a).
@@
A closer look shows, however, that the slowest rotators in our sample are all strongly boron depleted. In particular for the mid- and high-luminosity groups, our models can not reproduce all observed stars with the same mixing efficiency and age. While ESL 2 and 3 are marginally reproduced by our synthetic population, the nitrogen enhancement in ESL 2, (Hunter et al., 2009), implies that its boron abundance might be well below the determined upper limit. We consider this as an indication that a different boron depletion mechanism operated in the slow rotators, and postulate here, that the five stars with the smallest v sin i (ESL 16, 7, 10, 3 and 2) are intrinsic slow rotators, i.e., belong to the 25% of stars with a peculiar formation history, which may not be reflected by our single star models. While for each individual star this can not be proven, the likelihood that, e.g., ESL 16 belongs to the population of rapid rotators is less than 20%.
The remaining stars do show both expected trends, a stronger boron depletion for faster rotation and for higher luminosity, as predicted by our models, which can be exploited quantitatively. A close inspection of panels a-c of Fig. 7 shows that in all three luminosity ranges, less mixing than adopted in the stellar models would lead to a better agreement. This is even more so when an older age of NGC 3293 is considered. Panels d, e, and f in Fig. 7 illustrate the effect of reducing the efficiency parameter for rotational mixing by 50%, (). Notably, an older age can be compensated by a smaller mixing efficiency. Since an age of 12 Myr is at the lower end of the most recently proposed ages for NGC3293, (cf., Morel et al., 2022), we conclude that rotational mixing is likely less efficient than assumed in our models. Assuming we do interpret the small additional dip in the spectrum of ESL 05 (Fig. 5a, relative to the otherwise similar ESL 06 (Fig. 5b) as a detection of boron, then the resulting boron abundance of this star may provide the strongest constraint on a reduced rotational mixing efficiency.
@@
Perhaps, ESL 2 in our sample, and to a lesser extent also ESL 3, and 7, are the most interesting in this context. ESL 2 is an apparently slow rotator, strongly boron depleted, and the only star in our sample which is clearly nitrogen-enriched. Its location in the HR diagram clearly identifies it as a blue straggler (Fig. 2). Blue stragglers are thought to be the products of binary mass transfer or stellar mergers, (Schneider et al., 2015). According to recent models, merger products in the mass range considered here may be slow rotators (Schneider et al., 2019) and show various degrees of nitrogen enrichments, from none to strong, (Glebbeek et al., 2013). Wang et al. (2022) have also proposed that the slow rotators which form the bluer main sequence in young star clusters originate from stellar mergers. The boron abundances found in our sample stars may therefore fit into a consistent picture, with the majority of fast rotators enriching nitrogen and depleting boron by rotational mixing, and the slow rotators having a history of strong binary interaction. Slow rotators in which boron is not fully destroyed could support the suggestion of pre-main sequence mergers, (Tokovinin & Moe, 2020; Wang et al., 2022). To turn this idea from a working hypothesis into a theory, significantly more boron measurements, as well as merger models that include proton captures on boron, are required.